Calculate the linear expansivity or brass or length 170m,that assumes new length of 120.05m when heated through the temperature of 100⁰c

Calculate also the area and heated expansivity


1
Expert's answer
2022-03-16T18:29:22-0400

Solution.

"L_0=120m;"

"L=120.05m;"

"\Delta T=100^OC=100K;"

"L=L_0(1+\alpha_L t)\implies \alpha_L=\dfrac{L-L_0}{L_0\Delta T};"

"\alpha_L=\dfrac{120.05-120}{120\sdot100}=4.2\sdot 10^{-6}K^{-1};"

"\alpha_A=2\alpha_L;"

"\alpha_A=2\sdot4.2\sdot10^{-6}=8.4\sdot10^{-6}K^{-1};"

"\alpha_V=3\alpha_L;"

"\alpha_V=3\sdot4.2\sdot10^{-6}=12.6\sdot10^{-6}K^{-1};"

Answer: "\alpha_L=4.2\sdot 10^{-6}K^{-1};"

"\alpha_A=8.4\sdot10^{-6}K^{-1};"

"\alpha_V=12.6\sdot10^{-6}K^{-1}."